# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v2

@@ Appendix E Expert evaluation
We picked a random subset of $30$ generated notes and show them to $5$ general practitioners, in a random order. All evaluators got to see the same $30$ notes, together with the prompts that were used to generate them. All evaluators received detailed instructions on what was expected of them in the form of a PDF [^1] https://github.com/prabaey/SynSUM/blob/main/eval/Instructions˙clinical˙evaluation.pdf, which was orally explained by the researchers. The researchers then sat together with each evaluator separately to complete three example notes (different from the $30$ notes that were to be evaluated). Afterwards, the evaluators were asked to rate the notes in their own time, without the researchers’ involvement. Since it is infeasible to evaluate the whole dataset of $10{,}000$ notes, we opted for a small subset of $30$ notes, each going through a relatively extensive evaluation process that considered various measures of quality (evaluators took around $5-10$ minutes to evaluate each note). We decided to show all $5$ evaluators the same set of $30$ notes, to get a broader range of expert opinions in the evaluation of each note. This also allowed for the calculation of inter-annotator agreement.
@@
E.1 Evaluation aspects
Consistency
@@ E.1 Evaluation aspects
As explained in Section 2.2, around one third of the notes were generated using a second type of prompt, where the LLM is told that the patient does not suffer from any respiratory symptoms or underlying health conditions. Following the prompt in Figure 6, there are three parts of the prompt which can be violated (leading to penalties): (i) the respiratory symptoms, which the patient does not have, (ii) the underlying conditions, which the patient does not have, and (iii) the additional instructions. In our random set of $30$ notes, $20$ notes belonged to the first type, and $10$ to the second type.
@@
Quality of compact version
@@ E.2 Results
Extended results for consistency
@@
From this, we can conclude that a large majority of the inconsistencies arise from a violation of the additional instructions (usually by inventing additional symptoms), while the key symptom information included in the prompt was still conveyed correctly in the note. High consistency between the tabular variables and the concepts mentioned in the text makes our dataset a reliable resource for information extraction tasks.
Inter-Annotator Agreement
We calculate Krippendorff’s alpha with ordinal distance measure to get an idea of the inter-annotator agreement in our study. We get the following results:
@@
With an alpha of 1 indicating perfect agreement, we can say that there is some agreement between the 5 evaluators. Especially for consistency, readability and realism of the physical examination, the agreement is fair. Only the ratings for content of the compact version shows no indication of agreement. We hypothesize that this is at least partially due to the high scores assigned by all evaluators (mostly 4 and 5), with a deviation from 5 being assigned to random chance rather than to a common opinion of the evaluators. The lower standard deviations in Table 1 show that the evaluators consistently assign high scores to all aspects of the evaluation, even if they don’t agree on which particular notes deviate from these higher scores.
@@ Appendix F Symptom predictor baselines
Figure 9: Results for the neural-text baseline for the different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes. Figure best viewed in color.
@@ F.1 BN-tab
We learn a Bayesian network over the training data, providing the structure over all variables as in Figure 1. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library (Ankur Ankan and Abinash Panda 2015) with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2), and learn the parameters $p_{i}$ in Equations (3) through (6) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine built into pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for #days are learned using maximum likelihood estimation over the training set of $8{,}000$ examples. The likelihood is expressed as in Equation (7) and Equations (8) and (9) respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equations (7), (8) and (9) for all combinations of parent and child values. For the variable #days, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@
Once we have learned all parameters in the joint distribution, we can evaluate the Bayesian network’s ability to predict each of the symptoms. For each evidence setting (as defined in the main text), we apply variable elimination with each of the symptoms as a target variable. Looking at the causal structure in Figure 1, we note that the model never has to marginalize over the many rows in the learned #days CPT, since it is never a target variable. This makes automated inference feasible in our case.
@@ F.3 Neural-text
The pretrained BioLORD encoder (Remy, Demuynck, and Demeester 2024) was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys and mean all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
@@
Results for the different embedding types are reported numerically in Table 8, and shown visually in Figure 9. There is a significant gap in performance between the hist and phys settings for the symptoms cough, pain and fever. This makes sense, as the “history” section of the note outlines the symptoms experienced by the patient more clearly. The performance difference between the mean and concat settings is usually small.While the score for hist comes close to those for mean and concat, the latter usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions. For the compact notes, including the “physical examination” portion seems to confuse the model for the symptoms cough and pain, since the hist setting outperforms mean and concat there.
dyspnea cough pain nasal fever
normal
- hist 0.9399 0.9699 0.8310 0.9538 0.9028
- phys 0.9035 0.7948 0.6871 0.9526 0.8580
- mean 0.9660 0.9595 0.8415 0.9602 0.9125
- concat 0.9635 0.9660 0.8313 0.9606 0.9001
compact
- hist 0.9353 0.9557 0.8008 0.9522 0.8978
- phys 0.8747 0.7655 0.6492 0.9544 0.8324
- mean 0.9383 0.9480 0.7828 0.9583 0.9073
- concat 0.9358 0.9443 0.7849 0.9598 0.8964
Table 8: Results for the neural-text baseline using different embedding types at the input. We report F1-score over the test set for both the normal and compact versions of the notes. The best results per symptoms and per note version is shown in bold.

### v3

@@ Appendix C Additional example prompts
Appendix D Automated span extraction
To extract spans mentioning each symptom from the normal clinical notes, we send the following instructions to the LLM in the system message:
I will show you a clinical note containing information on a patient’s symptoms.
For each symptom, I will tell you whether the patient suffers from this symptom
or not.
Your task is to extract phrases from the note that mention these symptoms.
The annotation must have the following JSON structure:
      "symptom": one of the symptoms ("dyspnea", "cough", "respiratory pain",
      "fever" or "nasal symptoms")
      "text": phrase in the text that mentions the symptom and whether it is
      present or absent
      "symptom": ...
      "text":...
Keep the following instructions in mind:
- The same symptom may be mentioned multiple times. Include all phrases in
which a symptom is mentioned. Consider both the "history" portion of the note,
and the "physical examination" portion of the note.
- Also annotate a symptom if the note mentions that the patient does not
suffer from it.
- The phrases do not need to be full sentences, but need to be verbatim as
they appear in the note. You are not allowed to alter any words. If you
leave out words, use ...
- Order does not matter.
- You will reply only with the JSON itself, and you will not wrap in JSON
markers.
- You can only extract phrases from the "clinical note", not from any of
the other text in the prompt.
- Not all symptoms are necessarily mentioned in the note. Do not include a
symptom in the JSON if you cannot find any implicit or explicit mention of
it in the clinical note.
We then ask the LLM to extract the spans for each note, using the prompt shown in Fig. 4a.
For the compact notes, we tell the LLM to match the spans extracted from the normal note to the text in the compact note. We send the following instructions to the LLM in the system message:
I will show you two versions of a clinical note. The first version describes a
patient’s visit to the doctor’s office. The second one describes the same visit,
but in a more compact style (using abbreviations and shortcuts), while preserving
the overall message.
I will show you a set of phrases which were extracted from the first version of
the note. Your task is to map these to phrases in the second version of the note.
You must reply with the following JSON format.
   phrase in version 1 : corresponding phrase extracted in version 2
Keep the following instructions in mind:
- Please extract phrases verbatim.
- Please use the empty string if you cannot find a phrase with the same meaning.
- The phrases you extract must have the same meaning, you cannot simply copy
phrases that are in the same spot in the text.
- You will reply only with the JSON itself, and you will not wrap in JSON markers.
We then ask the LLM to match the phrases for each pair of notes, using the prompt shown in Fig. 4b. After the first response, we ask it to check if the extracted phrase has the same meaning, because we noticed that it often selected phrases because they were in the same spot in the text, even if they did not mean the same as the original phrase. We used the following prompt for this:
Please check if each extracted phrase has the same meaning as the original phrase.
If not, substitute it by the empty string (‘’). The rest of the JSON must remain
unchanged.
We post-process all phrases extracted from both the normal and compact notes by matching them to the text. If a phrase does not match the text exactly, we ask the LLM to correct itself. For this, we use the following prompt:
I will show you a clinical note, together with one or more phrases that were
extracted from it. However, some mistakes were made in extracting these
phrases. You must correct them.
The following is a clinical note:
< note >
The following phrases were extracted from this note. However, they do not
exactly match the text:
- < phrase 1 >
- < phrase 2 >
Please correct the phrases so they map exactly to a phrase in the text.
You must reply with the following JSON format:
   original phrase: corrected phrase
You will reply only with the JSON itself, and you will not wrap in JSON markers.
For all these prompts, we use OpenAI’s GPT-4o model. We set the temperature to 0.2 to encourage correctness over creativity. The full span annotation process cost around \$80. For a full walk through of the span extraction process, we refer to the tutorial notebook on our Github repository: https://github.com/prabaey/SimSUM/blob/main/src/span_annotations.ipynb.
@@ Appendix E Expert evaluation
We picked a random subset of $30$ generated notes and show them to $5$ general practitioners. All evaluators got to see the same $30$ notes, together with the prompts that were used to generate them, in a random order that differed between evaluators. All evaluators received detailed instructions on what was expected of them in the form of a PDF [^3] https://github.com/prabaey/SimSUM/blob/main/eval/Instructions_clinical_evaluation.pdf, which was orally explained by the authors. The authors then sat together with each evaluator separately to complete three example notes (different from the $30$ notes that were to be evaluated). Afterwards, the evaluators were asked to rate the notes in their own time, without the authors’ involvement. Since it is infeasible to evaluate the whole dataset of $10{,}000$ notes, we opted for a small subset of $30$ notes, each going through a relatively extensive evaluation process that considered various measures of quality (evaluators took around $5-10$ minutes to evaluate each note). We decided to show all $5$ evaluators the same set of $30$ notes, to get a broader range of expert opinions in the evaluation of each note. This also allowed for the calculation of inter-annotator agreement.
@@
E.1 Evaluation dimensions
E.1.1 Consistency
@@
As explained in Section 2.2, around one third of the notes were generated using a second type of prompt, where the LLM is told that the patient does not suffer from any respiratory symptoms or underlying health conditions. Following the prompt in Fig. 7, there are three parts of the prompt which can be violated (leading to penalties): (i) the respiratory symptoms, which the patient does not have, (ii) the underlying conditions, which the patient does not have, and (iii) the additional instructions. In our random set of $30$ notes, $20$ notes belonged to the first type, and $10$ to the second type.
@@ E.1.3 Clinical accuracy
E.1.4 Quality of compact version
@@ E.2 Results
E.2.1 Extended results for consistency
@@
E.2.2 Inter-Annotator Agreement
We calculate Krippendorff’s alpha with ordinal distance measure to quantify inter-annotator agreement:
@@ E.2.2 Inter-Annotator Agreement
With an alpha of 1 indicating perfect agreement, we can say that there is some agreement between the 5 evaluators. Especially for consistency, readability and realism of the physical examination, the agreement is fair. Only the ratings for content of the compact version shows no indication of agreement. We hypothesize that this is at least partially due to the high scores assigned by all evaluators (mostly 4 and 5), with a deviation from 5 being attributed to random chance rather than to a shared opinion of the evaluators. The lower standard deviations in Table 4 show that the evaluators consistently assign high scores to all aspects of the evaluation, even if they don’t agree on which particular notes deviate from these higher scores.
@@ Appendix F Symptom predictor baselines
Figure 10: Results for the neural-text baseline for the five different embedding types. We show F1-scores over the test set for both the normal and compact versions of the notes, averaging over 5 seeds. Figure best viewed in color. For numerical results, see Table 6 in Section 3.2.
@@ F.1 BN-tab
We learn a Bayesian network over the training data, providing the structure over all variables as in Fig. 2. For the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever and self-employed, we learn the conditional probability tables (CPTs) from the training data using maximum likelihood estimation (which comes down to counting co-occurrences of child and parent values for each entry in the CPT). We use the pgmpy library [1] with a K2 prior as a smoothing strategy to initialize empty CPTs.
@@
As support for learning Noisy-OR distributions is not provided in pgmpy, we learn these parameters with a custom training loop. We formulate the likelihood as in Equation (2) in Section 2.1, and learn the parameters $p_{i}$ in Equations (LABEL:eq:noisy_OR_dysp) through (LABEL:eq:noisy_OR_nasal) for the variables dyspnea, cough, pain and nasal through maximum likelihood estimation by iterating over the train set for 10 epochs, using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. To integrate the learned Noisy-OR distributions in the Bayesian network, we turn them into fully specified CPTs. To obtain these, we simply evaluate Equation (2) in Section 2.1 for all possible combinations of child and parent values. While this results in large and inefficient CPTs, the automated inference engine built into pgmpy library does not support Noisy-OR distributions directly. Note that both versions of the conditional distribution are equivalent, so we do not incur a loss in precision.
@@
Similarly, the coefficients in the logistic regression model for antibiotics and the Poisson regression model for #days are learned using maximum likelihood estimation over the training set of $8{,}000$ examples. The likelihood is expressed as in Equation (7), (8) and (9) in Section 2.1 respectively, with learnable parameters in place of each coefficient. We iterate over the train set for 15 epochs, again using an Adam optimizer with a batch size of 50, a learning rate of 0.01 and random initialization of each parameter. Finally, we turn the logistic regression and Poisson regression models into CPTs by evaluating Equation (7), (8) and (9) for all combinations of parent and child values. For the variable #days, we needed to turn each discrete number of days into a category, because pgmpy only provides automated inference for Bayesian networks consisting of exclusively categorical variables. This results in a large CPT containing one row per possible number of days, which range from 0 to 15 in our training dataset, and one column for each combination of the 7 parent variables. To allow for a possible larger maximum number of days in the test set, we create a category $\geq\text{{15 days}}$, which is defined as one minus the summed probability of all other days.
@@
Once we have learned all parameters in the joint distribution, we can evaluate the Bayesian network’s ability to predict each of the symptoms. For each evidence setting (as defined in the main text), we apply variable elimination with each of the symptoms as a target variable. Looking at the causal structure in Fig. 2, we note that the model never has to marginalize over the many rows in the learned #days CPT, since it is never a target variable. This makes automated inference feasible in our case.
@@ F.3 Neural-text
The pretrained BioLORD encoder [26] was obtained through the huggingface library. The encoder outputs 768-dimensional sentence embeddings. Since the full text did not fit into the context window, we embedded each sentence separately, and then combined them using our strategies outlined in the main text. To split the text into sentences, we used the nltk package. The settings hist, phys, mean and span all result in a text embedding of 768 dimensions, while the setting concat results in a text embedding of 2*768 dimensions.
@@
Results for the different embedding types are reported numerically in Table 6 in Section 3.2, and shown visually in Fig. 10.